OpenStudy (anonymous):

A company distributes free boxes of paperclips to all the workers in x offices. Each office has 8 more departments than the number of offices. The number of workers in each department is 4 less than the number of offices. Each worker is given 3 boxes of paperclips. Which expression shows the total number of boxes of paperclips each worker received? Answer choices: x x + 8 x + 4 3x(x + 8)(x + 4)

OpenStudy (anonymous):

@sylbot can you help me?

OpenStudy (anonymous):

The answer is the final choice. The reasoning goes something like this: x=number of offices. Each office has x+8 departments. Total # of departments is x(x+8). Each department has x-4 workers. Total workers= total # departments*(x-1)=x(x+8)(x-4) Each worker gets 3 boxes. Total box count is 3 times the total number of workers: 3x(x+8)(x-4)